OpenStudy (anonymous):

Solve for x. x-7/8 = 7/2 A. x=35 B. X=9(2/7) C.x= 31(1/2) D. x=63 (they are fractions)

OpenStudy (compassionate):

\[\frac{ x - 7 }{ 8} = \frac{ 7 }{ 2 }\] Is this it

OpenStudy (nincompoop):

always try to isolate on one side the variable you're trying to solve in order to achieve that, you need to figure out what are the parts that it comes with EXAMPLE: 3x-2b = 2f-3g if I want to solve for x in that example, I will need to make sure it is by itself on one side. I can achieve that by looking at the other parts like -2b. For me to get rid of -2b on the left side of the equation, I need to do something so it will turn into zero, that is add 2b since -2b + 2b is zero. Whatever I did on the left side must be done on the right side to keep the EQUALITY. So it will look like this 3x -2b + 2b = 2f -3g +2b it becomes 3x = 2f -3g + 2b now I need x not 3x. In order to achieve that I need to make sure that the 3 turns into 1 by dividing everything on the left-hand-side by 3 so I will be left with x. Again, what I do on the left side, must be done on the right side. 3x/3 = (2f -3g +2b)/3 it becomes x = (2f-3g+2b)/3 Try to do the same idea. Always isolate the variable. If you see something with an addition that needs to be "moved" on the other side of an equation, make sure you subtract; if you see a minus, make sure you add; if you see a multiplication, you divide; if you see a division, you multiple and so on and so fort.

OpenStudy (nincompoop):

HELLO? is it me you're looking for ???? pay attention next time when someone is trying to help you instead of leaving.

OpenStudy (anonymous):

\[\frac{ x-7 }{ 8 }=\frac{ 7 }{ 2 }\]First,multiply both sides by 8. Then,add 7 on both sides.\[(\frac{ x-7 }{ 8 })\times8=(\frac{ 7 }{ 2 })\times8\]\[x-7=28\]\[x-7+7=28+7\]\[x=?\]solve it @jojoheart